Differentiate the following with respect to x.

b) f(x) = ln 2x . sin3 (x2 – 3)


1
Expert's answer
2021-09-12T17:58:03-0400

Let us differentiate the function "f(x) = \ln 2x \cdot \sin 3 (x^2 – 3)" with respect to x using the product rule "(h(x)g(x))'=h'(x)g(x)+h(x)g'(x)" and the chain rule "h(g(x))'=h'(g(x))\cdot g'(x):"

"f'(x) = \frac{1}{2x}\cdot 2 \cdot \sin 3 (x^2 – 3)+\ln 2x \cdot \cos 3 (x^2 – 3)\cdot 3\cdot 2x\\
=\frac{ \sin 3 (x^2 – 3)}{x}+6x\cdot\ln 2x \cdot \cos 3 (x^2 – 3)."